OpenStudy (watermelon14):

If logx^2-log2a=log3a, then logx expressed in terms of loga is equivalent to i know the answer is (1/2)log6+loga but i don't get how

jhonyy9 (jhonyy9):

in this case do you know what is the index of logarithm when you write logx^2 for example ?

OpenStudy (watermelon14):

yeah 1/2logx

jhonyy9 (jhonyy9):

i tought it log_2 x so for example so this in case when you dont wrote there what number is in place of 2

OpenStudy (watermelon14):

i don't know what u mean

jhonyy9 (jhonyy9):

try this http://www.rapidtables.com/math/algebra/Logarithm.htm

OpenStudy (watermelon14):

yeah i am just confused about what you're asking

jhonyy9 (jhonyy9):

moment logx^2 -log2a = log3a add to both sides log2a what will get ?

jhonyy9 (jhonyy9):

logx^2 =log3a +log2a yes ?

OpenStudy (watermelon14):

yeah

jhonyy9 (jhonyy9):

logx^2 = 2logx bc. you know the formula log(a)^2 = 2log(a) yes ?

OpenStudy (watermelon14):

yeah

jhonyy9 (jhonyy9):

and so hope you know again that log a -log b = log (a/b)

jhonyy9 (jhonyy9):

do you can using in this above case ?

jhonyy9 (jhonyy9):

sorry there are log a +log b = log(a*b)

OpenStudy (mertsj):

\[\log(x^2)-\log(2a)=\log(3a) \]\[2\log(x)=\log(2a)+\log(3a)\] \[2\log(x)=\log(6a^2)\] \[2log(x)=\log(6)+\log(a^2)\] \[2\log(x)=\log(6)+2\log(a)\] \[\log(x)=\frac{1}{2}\log(6)+\log(a)\]

OpenStudy (watermelon14):

i see thanks you guys!!

jhonyy9 (jhonyy9):

np.